A 12 ȍ resistor has a potential difference of 4.0 volts. Determine the power dissipated in the resistor. 


1
Expert's answer
2021-12-24T11:32:48-0500

power dissipated in the resistor:


"P=UI=U^2\/R=4^2\/12=1.33" W